Josey drops a rock off a 100 meter cliff.
a) How long will it take for the rock to reach the ground?
b) What will be the impact velocity of the rock?

Answer by Boreal(15235)

s=(1/2)at^2
s(m)=(1/2)(9.8m^2/sec^2)*t^2 (sec^2)
200/9.8=t^2=20.41
t=sqrt(20.41)=4.52 seconds
v=sqrt(2gh)=sqrt(2*9.8 m/sec^2*100m) units are going to be m/sec
v=44.27 m/sec